# Measuring Visual Sycophancy in Multimodal Models

### v1

@@ Abstract
This paper introduces and examines the phenomenon of “visual sycophancy" in multimodal language models, a term we propose to describe these models’ tendency to disproportionately favor visually presented information, even when it contradicts their prior knowledge or responses. Our study employs a systematic methodology to investigate this phenomenon: we present models with images of multiple-choice questions, which they initially answer correctly, then expose the same model to versions with visually pre-marked options. Our findings reveal a significant shift in the models’ responses towards the pre-marked option despite their previous correct answers. Comprehensive evaluations demonstrate that visual sycophancy is a consistent and quantifiable behavior across various model architectures. Our findings highlight potential limitations in the reliability of these models when processing potentially misleading visual information, raising important questions about their application in critical decision-making contexts.
@@ 1 Introduction
Multimodal machine learning models, which integrate and reason across multiple modalities such as vision, language, and audio, have become increasingly significant in artificial intelligence research and applications (Manzoor et al., 2023; Zhang et al., 2023). These models develop richer representations by leveraging complementary information from diverse data types, enabling them to address complex tasks that span multiple domains (Anthropic, 2024; Achiam et al., 2023; Reid et al., 2024; Dubey et al., 2024).
@@
Vision-language models (VLMs), a subset of multimodal models, typically employ one of three main architectures: external vision encoders (Radford et al., 2021), cross-attention mechanisms (Alayrac et al., 2022; Tang et al., 2023), or end-to-end transformers (Anthropic, 2024; Achiam et al., 2023). While the specifics of these architectures differ, they all necessitate the integration of visual and textual modalities within the model to perform few-shot or zero-shot classification of provided images or to engage in natural conversation in relation to visual prompts
Modality Bias Distinct from Social Bias. It is necessary to distinguish the terminology of “bias” before we proceed with our investigation. Modality bias is distinct from the bias that the model displays due to the existence of unequal representation in the training data.
Social bias, a well-studied phenomenon in AI research, refers to the unfair or prejudiced treatment of certain groups or individuals based on characteristics such as race, gender, age, or socioeconomic status Ferrara (2023). In uni- and multimodal models, social bias often stems from imbalances or stereotypes present in the training data, leading to outputs that reflect and potentially amplify societal inequalities Starke et al. (2022); Akter et al. (2021); Shah and Sureja (2024). This type of bias has been extensively documented and remains a critical ethical concern in AI development.
However, our research focuses on a different, less explored form of bias: modality bias. Modality bias refers to the tendency of a multi-modal model to rely disproportionately on one modality (e.g., vision) over another (e.g., text) when making predictions or generating outputs Shtedritski et al. (2023); Guo et al. (2023). Unlike social bias, which primarily reflects societal inequalities from which the model samples, modality bias is a technical challenge inherent to the architecture and training of multi-modal systems. As foundational models are increasingly deployed to automate essential tasks across various domains, it is crucial to understand how training dynamics, architectural choices, causal mechanisms, and inference-time interventions may shape a model’s preference for one modality over another. This preference can lead to suboptimal performance or inconsistent behavior when processing multimodal inputs.
Bias as Systematic Deviation from the Mean. In probability theory and cognitive psychology, bias is often measured as a systematic additive or subtractive deviation from the mean, distinct from variance, which measures data spread. Probability Theory plus Noise (PT+N) model predicts systematic deviations from normative probabilities in human judgment, attributing these to random noise in memory sampling processes rather than additive biases in log odds (Howe and Costello, 2020). While the inner mechanisms of language models differ from human probability judgment, the concept of measuring bias as systematic deviation from the mean remains a useful observational metric in both domains.
Extending Keypoint Localization. Shtedritski et al. (2023) explored keypoint localization in vision-language models, particularly CLIP. By highlighting portions of the input image, they directed the model’s attention to specific subsets while maintaining global information, achieving state-of-the-art performance on referring expression comprehension tasks. Our work extends these localization methodologies to a broader range of large vision-language models, including both proprietary models and popular open-source models such as LLAVA (Liu et al., 2024). This comprehensive approach allows us to evaluate the generalizability of localization methods across diverse model architectures and training paradigms.
Modality Gap and The Platonic Representation Hypothesis. Recent research has investigated the modality gap or the separability of data points between modalities in the embedding space. Jiang et al. (2024) observed a significant gap between textual and visual information, finding that insufficient cross-modal representation alignment correlates with high rates of hallucinations. The Platonic Representation Hypothesis (PRH) posits that large-scale multimodal models may converge towards a shared statistical representation of reality Huh et al. (2024). While not the direct focus of our experimental work, this hypothesis provides context for understanding modality bias, suggesting that a model’s preference for one modality over another could indicate a divergence from a shared statistical representation.
Defining Visual Sycophancy and User Intent. Past research has examined various anthropomorphic, behavioral properties of large models, including corrigibility Perez et al. (2022), sycophancy Sharma et al. (2023); Denison et al. (2024), and truthfulness Campbell et al. (2023); Evans et al. (2021). Our study focuses on visual sycophancy, a specific case where a concept or object is redundantly presented across two modalities (text and vision), and the model displays a systematic bias towards the option that shows user intent or emphasis.
We examine this phenomenon by using keypoint localization as a proxy for user intent, where highlighted portions of the image or contrastively colored texts provide additional information not present in the text alone Shtedritski et al. (2023). Through counterfactual comparisons between neutral conditions and biased variations, we analyze how visual cues affect the model’s responses, measuring both the direction and magnitude of changes. This approach allows us to assess the extent to which the model may prioritize perceived user preferences over providing the most accurate or truthful response.
@@
Building upon the theoretical framework of modality bias and visual sycophancy outlined in our introduction, we now turn to the practical challenge of empirically measuring these phenomena in large language models. Our methodology aims to operationalize the concept of visual sycophancy through a series of carefully designed experiments. By transforming established benchmarks into multimodal representations and systematically manipulating visual cues, we create a controlled environment to observe how models respond to subtle visual emphasis. This approach allows us to quantify the extent to which models prioritize visually emphasized information over their prior knowledge or training, directly addressing our research questions about the nature and extent of visual sycophancy in multimodal AI systems. In the following section, we detail our experimental design, data preparation, and analytical methods, providing a clear path from our theoretical considerations to concrete, measurable outcomes.
@@ 2 Method
To empirically investigate the phenomenon of visual sycophancy in multimodal language models, we designed a series of experiments to quantify the influence of visual cues on model outputs.
Evaluation Bias. Quantifying a model’s bias on a multiple-choice benchmark is relatively straightforward compared to quantifying bias on a free-generation task. Nevertheless, quantification remains challenging due to the multidimensional nature of bias. In the context of social biases, a model’s output may be biased in content, style, or framing Bang et al. (2024), necessitating multidimensional metrics. However, in the context of modality bias that skews towards a localized key point, where the evaluation metric is percent correctness and a ground truth answer exists, a one-dimensional metric may effectively capture and quantify the bias, as the focus is on a single, well-defined aspect of performance.
@@
Our evaluation of bias measures the shift in the distribution of answers and log probabilities between biased variations. This involves comparing answer distributions across variations, including the distribution of ground truth answers. We also calculate a bias percentage, which represents the proportion of answers that changed in both directions. For GPT-4o-mini and LLAVA, we conduct an analysis of the shift in the distribution of the top 4 answer token log probabilities across variations.
@@
Pre-marked Visual Prompt Generation. For each neutral prompt in our benchmarks, we generated biased variations through visual localization techniques. This process involved creating visually distinct versions of the same question, each emphasizing a different answer option to simulate user intent.
For the Visual MMLU (vMMLU) benchmark, we produced two types of variation formats. The first format uses a filled-in bubble with colored text: we filled in the bubble next to one answer option and colored its text, creating a visual emphasis on that option. This highlight suggests user intent. The second format employs a size variation: we doubled the font size of the biased option without highlighting or bubbling the option. This size change does not explicitly suggest user intent.
@@
For the Visual Social IQa (vSocialIQa) benchmark, we also created two types of formats. The first format uses bubbled and highlighted text: we filled in the bubble next to one answer option and highlighted its text in yellow. The second format adopts a web-format style: we designed a variation resembling a typical web format, featuring a light background, black text for the question and options, and a light blue highlight for one answer option.
@@
Visual Prompt Dimension. To ensure consistency across our experiments, we maintained uniform input dimensions for all prompts and their variations within each benchmark type. This standardization was crucial for a fair comparison across different variations and models. The specific dimensions were chosen to accommodate the visual elements of our biased variations while preserving the readability and structure of the original prompts. For vMMLU, the image prompts were consistently 560 x 640 pixels, and for vSocialIQa, the image prompts were 800 x 600 pixels.
@@
vMMLU and vSocialIQa. MMLU primarily assesses factual knowledge and mathematical computation skills, while Social IQa evaluates a model’s capacity to provide socially appropriate responses in various situations Sap et al. (2019); Hendrycks et al. (2021). Both the vMMLU and vSocialIQa variations measure whether there is a significant correlation between shifts in answer distribution and the visual bias presented by keypoint localization. We chose to experiment with visual sycophancy on both general knowledge and social intelligence tasks to disentangle inconsistencies in factual and social contexts. This approach allows us to examine whether the effects of visual sycophancy differ between tasks that require objective information and those that involve nuanced social reasoning.
@@ 3 Result
Our experiments on the vMMLU benchmark reveal patterns of visual sycophancy across the tested multimodal language models, with varying degrees of susceptibility among different architectures. Table 1 summarizes these findings.
@@ A B C D (%) (%) (%)
LLAVA and Claude-haiku demonstrated the most resilience to visual sycophancy. We observed that LLAVA’s performance on vMMLU is at around 50 percent in the neutral setting and does not undergo significant change for pre-marked options. After conducting paired statistical t-tests to compare the distribution shifts between the neutral setting and each of the biased conditions (Options A, B, C, and D), the results indicated that the shifts in answer distributions are not statistically significant. Claude-haiku also demonstrated resilience to visual sycophancy, with relatively modest shifts in response distribution. Its largest shift occurred when Option C was pre-marked, resulting in a 20 percentage point increase. Interestingly, Claude-haiku’s overall performance remained relatively stable across different pre-marking conditions, with score changes ranging from 0 to -7 percentage points. While Claude-haiku is not immune to visual sycophancy, its impact is limited.
@@ A B C (%) (%) (%) A B C (%) (%) (%)
In contrast, Gemini-1.5-flash exhibited the highest susceptibility to visual sycophancy. It showed substantial shifts in response distribution for all pre-marked options, with increases ranging from 30.06 to 44.98 percentage points. These large shifts were accompanied by more pronounced decreases in non-marked option selection, averaging between -10.02 and -14.99 percentage points. Notably, Gemini-1.5-flash also experienced the most significant performance degradation, with score decreases of up to 23.18 percentage points when Option C was pre-marked. The bias was so pronounced that visual variations could be classified based on the responses alone, with Gemini selecting visually highlighted options A or B about 50 percent of the time and options C or D approximately 70 percent of the time. This pattern suggests a strong influence of visual cues on Gemini-1.5-flash’s decision-making at the expense of accuracy. Interestingly, Gemini demonstrated a notable difference in its behavior between blue-centered and vanilla vSocialIQa formats, indicating that the visual presentation of options can significantly impact model responses, potentially mitigating or exacerbating biases.
@@
GPT-4o-mini displayed an intermediate level of susceptibility to visual sycophancy. Its shifts towards pre-marked options, while noticeable, were generally less pronounced than those of Gemini-1.5-flash but more substantial than Claude-haiku’s. Interestingly, GPT-4o-mini’s performance impact varied depending on the pre-marked option, with both slight improvements and decreases observed. This variability hints at a complex interaction between visual cues and the model’s underlying knowledge or decision-making processes.
@@
The strength of the sycophancy effect also varied depending on which option was pre-marked. Across all models, pre-marking Options C and D generally elicited stronger effects compared to Options A and B. This pattern raises questions about potential positional biases or the influence of option ordering on the models’ susceptibility to the visual cue. The observations from the vMMLU benchmark are further nuanced by the results from the vSocialIQa task, as shown in Table 2. The vSocialIQa results not only corroborate the presence of visual sycophancy across different task types but also reveal how the effect can be modulated by subtle changes in visual presentation.
@@
Visual design significantly influences sycophancy effects. The comparison between Setup A and Setup B in the vSocialIQa task demonstrates that the visual presentation of options can dramatically alter the magnitude of visual sycophancy. This is most strikingly illustrated by Claude-haiku’s performance. While it showed resilience in the vMMLU task and in vSocialIQa Setup A, it exhibited extreme susceptibility in Setup B, with shifts of up to 65 percentage points when Option B was pre-marked and a complete 100% selection of Option C when it was pre-marked. This stark contrast underscores the critical role of visual design in multimodal tasks and suggests that model behavior can be highly sensitive to seemingly minor changes in presentation.
@@
Consistency of susceptibility varies across models. Gemini-1.5-flash’s high susceptibility to visual sycophancy, observed in the vMMLU task, is consistently evident across both vSocialIQa setups. This reinforces the notion that some models may have a more fundamental vulnerability to visual cues, regardless of the specific task or visual presentation. In contrast, GPT-4o-mini’s intermediate level of susceptibility in vMMLU is mirrored in its relatively stable behavior across both vSocialIQa setups, suggesting a more robust integration of visual and textual information.
@@ 4 Analysis
Changes in Token Probability. We begin our analysis by examining the shift in log probability. Log probabilities for the answer tokens (‘A,’ ‘B,’ ‘C,’ and ‘D’) were collected shortly after inference. For each visual bias type, the change in token probability from neutral to bias were calculated and then averaged out across prompts. The probabilities were first converted back to linear probability to avoid unintended scaling of the values during calculation of deltas.
@@
As Figure 5 shows, the bias type strongly aligns with the answer choice, with the highest positive delta from neutral to bias. That is, the model is most likely to increase the probability for token X when the visual information suggests X. Although the distribution of the model’s answers may capture a model’s tendency to prefer one answer over another, it has been observed that first-token probability does not always reflect the model’s “actual” answer during generation (Wang et al., 2024). Thus, an in-depth analysis of the delta in token probability was necessary.
Figure 6 displays the changes in token probability for LLAVA-1.5v-13b, which shows a different type of bias. Instead of showing a stronger preference for the biased option as GPT-4o-mini did, LLAVA showed a stronger preference for option D across all biased variations, with the exception of variation A, in which options C and D were tied for positive delta.

### v2

# Measuring Agreeableness Bias in Multimodal Models

@@ Abstract
This paper examines a phenomenon in multimodal language models where pre-marked options in question images can significantly influence model responses. Our study employs a systematic methodology to investigate this effect: we present models with images of multiple-choice questions, which they initially answer correctly, then expose the same model to versions with pre-marked options. Our findings reveal a significant shift in the models’ responses towards the pre-marked option, even when it contradicts their answers in the neutral settings. Comprehensive evaluations demonstrate that this agreeableness bias is a consistent and quantifiable behavior across various model architectures. These results show potential limitations in the reliability of these models when processing images with pre-marked options, raising important questions about their application in critical decision-making contexts where such visual cues might be present. We share our research code at github.com/jasonlim131/looksRdeceiving.
@@ 1 Introduction
Multimodal language models, which integrate and reason across multiple modalities such as vision, language, and audio, have become increasingly significant in artificial intelligence research and applications (Manzoor et al., 2023; Zhang et al., 2023). These models develop richer representations by leveraging complementary information from diverse data types, enabling them to address complex tasks that span multiple domains (Anthropic, 2024; Achiam et al., 2023; Reid et al., 2024; Dubey et al., 2024). Vision-language models (VLMs), a subset of multimodal models, typically employ one of three main architectures: external vision encoders (Radford et al., 2021), cross-attention mechanisms (Alayrac et al., 2022; Tang et al., 2023), or end-to-end transformers (Anthropic, 2024; Achiam et al., 2023). While the specifics of these architectures differ, they all necessitate the integration of visual and textual modalities within the model to perform few-shot or zero-shot classification of provided images or to engage in natural conversation in relation to visual prompts.
Our research focuses on a specific phenomenon in multimodal models: their tendency to change their answers when presented with visually pre-marked options, even when these options contradict their prior knowledge or initial responses. This effect, which we refer to as “agreeableness bias,” raises important questions about the reliability and consistency of multimodal model outputs in scenarios where visual cues might influence decision-making. To systematically investigate this phenomenon, we designed a series of experiments using established benchmarks transformed into visual representations. Our methodology involves presenting models with images of multiple-choice questions, first in a neutral format and then in versions with pre-marked options. By comparing the models’ responses across these variations, we can quantify the extent to which visual pre-marking influences their decision-making.
We examine this effect across different types of tasks, including the Visual MMLU (vMMLU) benchmark, which assesses factual knowledge and mathematical computation skills, and the Visual Social IQa (vSocialIQa) benchmark, which evaluates a model’s capacity to provide socially appropriate responses in various situations Sap et al. (2019); Hendrycks et al. (2021). This approach allows us to explore whether the agreeableness bias differs between tasks requiring objective information and those involving nuanced social reasoning.
Our study encompasses a range of multimodal models, including both proprietary and open-source architectures such as Claude-haiku, Gemini-1.5-flash, GPT-4o-mini, and LLAVA. By analyzing the behavior of these diverse models, we aim to understand the prevalence and variability of agreeableness bias across different model designs and training paradigms. The findings from our experiments reveal significant shifts in model responses towards pre-marked options, with the magnitude of these shifts varying across models and task types. These results highlight potential limitations in the reliability of multimodal models when processing images with pre-marked options, raising important considerations for their application in critical decision-making contexts.
In the following sections, we detail our experimental design, present our findings, and discuss their implications for the development and deployment of multimodal AI systems. Our work contributes to the ongoing discussion about the robustness and reliability of AI models, particularly in scenarios where visual cues may influence their outputs.
@@ 2 Method
To empirically investigate the phenomenon of agreeableness bias in multimodal language models, we designed a series of experiments to quantify the influence of pre-marked options on model outputs.
@@
Evaluation of Bias. We measured the shift in the distribution of answers and log probabilities between neutral and biased variations. This involved comparing answer distributions across variations, including the distribution of ground truth answers. We calculated a bias percentage, representing the proportion of answers that changed in both directions. For GPT-4o-mini and LLAVA, we analyzed the shift in the distribution of the top 4 answer token log probabilities across variations.
@@
Pre-marked Visual Prompt Generation. For each neutral prompt in our benchmarks, we generated biased variations by visually pre-marking different answer options. This process created visually distinct versions of the same question, each emphasizing a different answer option.
For the Visual MMLU (vMMLU) benchmark, we produced two types of variation formats:
(1) Filled-in bubble with colored text: We filled in the bubble next to one answer option and colored its text, creating a visual emphasis on that option.
(2) Size variation: We doubled the font size of the biased option without highlighting or bubbling the option.
For the Visual Social IQa (vSocialIQa) benchmark, we also created two types of formats:
(1) Bubbled and highlighted text: We filled in the bubble next to one answer option and highlighted its text in yellow.
@@
(2) Web-format style: We designed a variation resembling a typical web format, featuring a light background, black text for the question and options, and a light blue highlight for one answer option.
@@
Visual Prompt Dimensions. We maintained uniform input dimensions for all prompts and their variations within each benchmark type. For vMMLU, the image prompts were consistently 560 x 640 pixels, and for vSocialIQa, the image prompts were 800 x 600 pixels.
@@
vMMLU and vSocialIQa Benchmarks. We chose these two benchmarks to examine whether the effects of agreeableness bias differ between tasks that require objective information (MMLU) and those that involve nuanced social reasoning (Social IQa) Sap et al. (2019); Hendrycks et al. (2021). Both variations measure whether there is a significant correlation between shifts in answer distribution and the visually pre-marked options. We chose to experiment with agreeableness bias on both general knowledge and social intelligence tasks to disentangle inconsistencies in factual and social contexts. This approach allows us to examine whether the effects of agreeableness bias differ between tasks that require objective information and those that involve nuanced social reasoning.
Our methodology allows us to quantify the extent to which pre-marked options influence model responses, potentially overriding their prior knowledge or training. By comparing model behaviors across different visual presentations and task types, we aim to provide a comprehensive understanding of agreeableness bias in multimodal AI systems.
@@ 3 Result
Our experiments on the vMMLU and vSocialIQa benchmarks reveal patterns of agreeableness bias across the tested multimodal language models, with varying degrees of susceptibility among different architectures. Table 1 summarizes the findings for the vMMLU benchmark.
@@ A B C D (%) (%) (%)
LLAVA and Claude-haiku demonstrated the most resilience to agreeableness bias We observed that LLAVA’s performance on vMMLU is at around 50 percent in the neutral setting and does not undergo significant change for pre-marked options. Claude-haiku also demonstrated resilience to agreeableness bias, with relatively modest shifts in response distribution. Its largest shift occurred when Option C was pre-marked, resulting in a 20 percentage point increase. Interestingly, Claude-haiku’s overall performance remained relatively stable across different pre-marking conditions, with score changes ranging from 0 to -7 percentage points. While Claude-haiku is not immune to agreeableness bias, its impact is limited.
@@ A B C (%) (%) (%) A B C (%) (%) (%)
In contrast, Gemini-1.5-flash exhibited the highest susceptibility to agreeableness bias. It showed substantial shifts in response distribution for all pre-marked options, with increases ranging from 30.06 to 44.98 percentage points. These large shifts were accompanied by more pronounced decreases in non-marked option selection, averaging between -10.02 and -14.99 percentage points. Notably, Gemini-1.5-flash also experienced the most significant performance degradation, with score decreases of up to 23.18 percentage points when Option C was pre-marked. The bias was so pronounced that visual variations could be classified based on the responses alone, with Gemini selecting visually highlighted options A or B about 50 percent of the time and options C or D approximately 70 percent of the time. This pattern suggests a strong influence of visual cues on Gemini-1.5-flash’s decision-making at the expense of accuracy. Interestingly, Gemini demonstrated a notable difference in its behavior between blue-centered and vanilla vSocialIQa formats, indicating that the visual presentation of options can significantly impact model responses, potentially mitigating or exacerbating biases.
@@
GPT-4o-mini displayed an intermediate level of susceptibility to agreeableness bias. Its shifts towards pre-marked options, while noticeable, were generally less pronounced than those of Gemini-1.5-flash but more substantial than Claude-haiku’s. Interestingly, GPT-4o-mini’s performance impact varied depending on the pre-marked option, with both slight improvements and decreases observed. This variability hints at a complex interaction between visual cues and the model’s underlying knowledge or decision-making processes.
@@
The strength of the agreeableness bias also varied depending on which option was pre-marked. Across all models, pre-marking Options C and D generally elicited stronger effects compared to Options A and B. This pattern raises questions about potential positional biases or the influence of option ordering on the models’ susceptibility to the visual cue. The observations from the vMMLU benchmark are further nuanced by the results from the vSocialIQa task, as shown in Table 2. The vSocialIQa results not only corroborate the presence of agreeableness bias across different task types but also reveal how the effect can be modulated by subtle changes in visual presentation.
@@
Visual design modulates agreeableness effects. The comparison between Setup A and Setup B in the vSocialIQa task demonstrates that the visual presentation of options can dramatically alter the magnitude of agreeableness bias. This is most strikingly illustrated by Claude-haiku’s performance. While it showed resilience in the vMMLU task and in vSocialIQa Setup A, it exhibited extreme susceptibility in Setup B, with shifts of up to 65 percentage points when Option B was pre-marked and a complete 100% selection of Option C when it was pre-marked. This stark contrast underscores the critical role of visual design in multimodal tasks and suggests that model behavior can be highly sensitive to seemingly minor changes in presentation.
@@
Consistency of susceptibility varies across models. Gemini-1.5-flash’s high susceptibility to agreeableness bias, observed in the vMMLU task, is consistently evident across both vSocialIQa setups. This reinforces the notion that some models may have a more fundamental vulnerability to visual cues, regardless of the specific task or visual presentation. In contrast, GPT-4o-mini’s intermediate level of susceptibility in vMMLU is mirrored in its relatively stable behavior across both vSocialIQa setups, suggesting a more robust integration of visual and textual information.
@@ 4 Analysis
Our analysis delves deeper into the patterns of agreeableness bias observed in the results, focusing on changes in token probability and the varying susceptibility of different models across tasks and visual formats.
Changes in Token Probability We examined the shift in log probability for answer tokens (’A,’ ’B,’ ’C,’ and ’D’) across different visual bias conditions. For each bias type, we calculated the change in token probability from neutral to biased conditions, averaged across prompts. To avoid unintended scaling effects, we first converted the probabilities back to linear probability before calculating deltas. As shown in Figure 5, for GPT-4o-mini, the bias type strongly correlates with increased token probability for the corresponding answer choice. This indicates that the model is most likely to increase the probability of a given token when the visual information suggests that the token is the answer. Interestingly, LLAVA-1.5v-13b exhibited a different pattern of bias, as shown in Figure 6. Instead of showing a stronger preference for the biased option, LLAVA consistently showed a stronger preference for option D across all biased variations, with the exception of variation A, where options C and D were tied for positive delta.